OpenStudy (chunkymonkay):

Will Give Medal :) Two cars traveled equal distances in different amounts of time. Car A traveled the distance in 4 h, and Car B traveled the distance in 4.5h. Car B traveled 5 mph slower than car A. How fast did Car B travel?

OpenStudy (nurali):

x-5=speed of car B x = speed of car A Distance = speed times time distances is equal so (x-5)(4.5)=(x)(4) 4.5x-22.5=4x multiply both sides by 2 to get rid of the decimals 9x-45=8x subtract 8x from both sides x-45=0 a add 45 to both sides and get x=45 speed of car A = 45 mph Speed of car B = 45-5 or 40 mph Speed of car B = 40 mph

OpenStudy (chunkymonkay):

Thank you!!

OpenStudy (nurali):

My Pleasure